# AsynFusion: Towards Asynchronous Latent Consistency Models for Decoupled Whole-Body Audio-Driven Avatars

### v1

@@ 1 Introduction
Audio-driven avatar expression and pose generation [8, 56, 29, 2, 5, 29, 30] is a crucial task aimed at creating lifelike digital humans that can seamlessly translate audio input into synchronized facial expressions and body poses. This task is fundamental to bridging the gap between speech and nonverbal communication, enabling avatars to convey emotions, intentions, and personality in a natural and dynamic manner. Its importance spans a wide range of fields, including metaverse applications, digital human development, gaming, and human-computer interaction [19, 55, 41, 57].
In recent years, numerous methods have been proposed for audio-driven avatar expression and pose generation, primarily treating speech-driven facial expression and body motion synthesis as separate tasks. Facial expression generation [13, 45] focuses on mapping emotional features from speech to facial muscle movements for natural animations, while body motion synthesis [18, 10] explores correlations between speech and gestures to generate coherent full-body motions. Despite advancements, these methods often lack sufficient coordination between expressions and movements. Generative models like VQ-VAE [2], GANs [20], and diffusion models [1, 3, 9, 58, 49, 44] have improved synchronization and diversity, enabling unified modeling of expressions and movements [20, 9, 37, 33]. As shown in Fig. 1, recent works include Probtalk [37], which generates expressions and postures simultaneously with a unified model, DiffSHEG [9], which uses a unidirectional sequence from expression to gestures, and EMAGE [33], which incorporates body hints for better coordination. Combo [60], the most related work, combines features for expressions and movements into a joint bidirectional distribution. However, a key challenge remains: balancing coordination accuracy and computational efficiency. More specifically, synchronization of expressions and movements often incurs high computational overhead, limiting the production of fluid animations in latency-sensitive scenarios.
To address this challenge, we propose AsynFusion, a novel framework that leverages the decoupling of facial expression and body posture generation to achieve efficient and lifelike animations. The core idea of AsynFusion lies in separating the generation processes of the head and body to facilitate parallel information processing while ensuring shared feature interactions. This design not only respects the different information requirements of facial and gestural elements but also incorporates asynchronous mechanisms to optimize computational efficiency without compromising coordination quality. Building upon this principle, we design a model with three key components: (1) A dual-branch Diffusion Transformer architecture that enables parallel generation of facial expressions and body gestures, ensuring their alignment through bidirectional feature interactions. (2) A cooperative synchronization module that employs cross-attention mechanisms to model the intricate dependencies between facial and gestural elements, enhancing the coherence and natural dynamics of the generated animations. (3) An Asynchronous Latent Consistency Model(LCM) Sampling method that reduces computational complexity while maintaining high-quality results, enabling real-time performance suitable for interactive applications.
@@ 2.1 Speech-driven Expression Generation
Research in speech-driven facial animation has seen significant advancement over the past few decades. Early approaches were predominantly rule-based [14, 28, 29], offering explicit control over facial animations but often requiring labor-intensive manual intervention to achieve high-quality results. With the rise of deep learning, the field shifted towards data-driven methods [7, 12, 26, 27, 35, 36] which aimed to generate more natural speech-synchronized facial animations. While these approaches achieved promising results, particularly in image generation, they frequently encountered challenges such as pixel-level artifacts and inconsistencies in 3D geometry. Recent progress in 3D facial animation [27] has addressed some of these limitations, with notable advances driven by transformer-based architectures such as FaceFormer [16] and CodeTalker [59]. These methods introduced specialized attention mechanisms that improved temporal alignment between speech and facial expressions. Despite these improvements, most current models adopt deterministic mappings, which limit the diversity of facial expressions and eye movements, even though lip movements are closely aligned with speech. Our work builds upon these advancements by addressing the limitations of deterministic mappings and enhancing the expressiveness of facial animations.
@@ 2.2 Speech-driven Gesture Generation
The evolution of gesture generation closely mirrors the advancements in facial animation, transitioning from early rule-based systems [8, 56, 29] to more sophisticated data-driven approaches [2, 5, 29, 30]. Initial data-driven methods primarily utilized statistical models to map speech to gestures, and were later followed by deterministic architectures such as MLPs [30], CNNs [20], RNNs [5, 21, 34], and more recently, Transformers [6]. Recognizing the many-to-many relationship between speech and gesture, where a single speech input can correspond to a diverse range of gestures, recent methods have embraced generative frameworks. These include GANs [20], and diffusion models [1, 9, 61, 58], which offer greater flexibility and diversity in gesture synthesis. Among diffusion-based approaches, DiffGesture [68] and DiffuseStyleGesture [62] were pioneering works that applied diffusion models to gesture generation. However, these methods typically require initial motion seeds to generate longer sequences, which can limit their applicability in continuous gesture generation. LDA [1] introduced translation-invariant positional embeddings to address longer sequence generation, but it still encounters challenges when dealing with streaming inputs or very long sequences. Building on these insights, our work aims to overcome the limitations of existing diffusion-based methods by introducing a more robust framework for continuous gesture generation, without relying on predefined motion seeds.
@@ 2.3 Joint Expression-Gesture Generation
Recent efforts have started to explore the joint generation of facial expressions and body gestures, aiming to create more synchronized and lifelike animations. Early work by Habibie et al. [20] introduced a CNN-based framework that used shared speech encoding and multiple decoders for different body parts, enhanced by adversarial training. However, their deterministic approach struggled to capture the natural variability in the mapping between speech and motion, leading to less dynamic and diverse outputs. Yi et al. [65] proposed a more modular method, using Wav2Vec [4] to drive facial expressions and VQ-VAE for gesture generation. While this token-based approach improved task separation, it also constrained motion diversity, as the discretization limited the range of generated gestures. [37] employs a unified model to learn the mapping from audio to both expressions and poses. EMAGE [33] explored incorporating body hints to enhance joint generation outcomes. A notable advancement came with DiffSHEG [9], which introduced a unified diffusion-based framework where information flowed unidirectionally from facial expressions to body gestures. This approach improved coordination between facial and body movements, but the unidirectional design restricted the rich, mutual influence that naturally occurs between expressions and gestures. Although Combo [60] attempts to model the bidirectional feature distribution between facial expressions and poses, its synchronous generation design hinders more efficient results. In our work, we aim to address these limitations by enabling more flexible interaction between facial expressions and gestures, while also improving the efficiency of the generation process. Our method promotes better coordination without sacrificing the diversity and natural variability of the generated motions.
@@ 3.1 Preliminary
3.1.1 Diffusion Transformers
Diffusion Transformers (DiT) employ a latent diffusion model (LDM) with a Transformer backbone for motion generation. Let $\mathbf{x}_{0}^{E}$, $\mathbf{x}_{0}^{G}$, and $\mathbf{x}_{0}^{M}$ denote expressions, gestures, and motion clips, respectively, aiming to model the motion distribution $p(\mathbf{x}_{0})\in\mathbb{R}^{N\times(3J+D_{exp})}$, where $N$ is the number of frames, $J$ the skeletal joints, and $D_{exp}$ the expression blend shape dimension.
@@ 3.1.1 Diffusion Transformers
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$
@@
$q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$
@@
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$
@@
$\epsilon_{\theta}(\mathbf{x}_{t},t,c)=f_{i}\circ f_{i-1}\circ\dots\circ f_{1}(\mathbf{x}_{t},t,c),$
@@
3.1.2 Latent Consistency Models
@@ 3.2 The AsynFuison Framework
Dual-branch DiT Architecture. To address the challenges of generating coordinated facial expressions and body gestures, AsynFusion introduces a Dual-branch DiT Architecture. This architecture consists of two parallel branches, each designed to specialize in its respective motion domain: the expression branch focuses on capturing subtle facial movements and lip synchronization, while the gesture branch handles broader body dynamics and pose transitions. By using separate transformer blocks with independent parameters, the branches are able to learn domain-specific temporal dependencies, ensuring highly specialized motion generation for each modality. For each branch, the input comprises three key components: (1) noisy motion samples $\mathbf{z}_{t}^{E}$ or $\mathbf{z}_{t}^{G}$, generated by applying Gaussian noise to the target expressions or gestures, (2) timestep embeddings $\gamma(t)$, encoding the noise level at time step $t$, and (3) audio features $\mathbf{F}_{\mathrm{aud}}$, which serve as conditioning signals for both branches to ensure synchronized generation:
$\hat{\mathbf{z}}^{E}$ $=\mathcal{T}_{E}(\mathbf{z}_{t}^{E},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$ (4)
$\hat{\mathbf{z}}^{G}$ $=\mathcal{T}_{G}(\mathbf{z}_{t}^{G},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$
where $\mathcal{T}_{E}$ and $\mathcal{T}_{G}$ represent the expression and gesture transformers, respectively. The shared audio features enable each branch to independently map audio to motion while maintaining temporal alignment. This design allows the branches to specialize in their respective motion types, with cross-branch interaction through a synchronization mechanism. Additionally, the branch separation supports asynchronous sampling during inference, adapting to the different temporal dynamics of expressions and gestures.
Cooperative Synchronization Module. To orchestrate the intricate interplay between facial expressions and body gestures, we propose the Cooperative Synchronization (CoSync) module as an integral component that facilitates sophisticated bidirectional interaction between the dual branches of our architecture. After each Transformer block, a bi-Flow cooperative synchronization layer built upon the cross-attention mechanism is introduced to preliminarily model the relationship between the two branches, capturing holistic dynamic priors to enhance the performance of each component. CoSync module creating a dynamic neural bridge that facilitates seamless information exchange between expression and gesture representations. We take gesture to expression data-flow for example, the query $\mathbf{Q}_{exp}$ is extracted by linear projection from $\mathbf{F}_{i}^{exp}$ ($i$ is the layer index), and the key and value $\mathbf{K}_{ges}$, $\mathbf{V}_{ges}$ are extracted from $\mathbf{F}_{i}^{ges}$ in the same way. To obtain the updated facial feature $\tilde{\mathbf{F}}_{i}^{exp}$,
@@
where MLP and LN is a MLP block and a LayerNorm, $\sqrt{d}$ is a scaling factor. This bidirectional feature exchange enables the model to capture subtle correlations between facial micro-expressions and corresponding gestural nuances, much like the natural synchronization observed in human behavior. What distinguishes the CoSync module is its ability to maintain the delicate balance between modality-specific independence and cross-modal coordination. While each branch preserves its specialized focus, the module enables them to share complementary information that enhances the overall coherence of the generated animation. This sophisticated interaction mechanism ensures that facial expressions and body gestures achieve natural synchronization without enforcing rigid correlations, resulting in more fluid and lifelike avatar animations.
@@
where the expression branch typically requires fewer sampling steps ($T_{exp}$) compared to the gesture branch ($T_{ges}$) due to its more constrained motion space. To maintain the bidirectional feature interaction during asynchronous sampling, we introduce a dynamic feature buffer mechanism within the CoSync module. At each sampling step, intermediate features from both branches are stored in the buffer and updated asynchronously. When one branch performs a denoising step, it can still access the most recent features from the other branch through the buffer, enabling continuous cross-modal information exchange even with different sampling rates. This asynchronous sampling strategy ensures that the expression branch can maintain rapid generation for time-critical facial movements while incorporating relevant gestural context, the gesture branch can use more sampling steps to capture complex body dynamics while staying informed by facial expressions, and the bidirectional feature interaction by our CoSync module remains active throughout the entire generation process, preserving motion coherence despite different sampling frequencies which resulting in both efficient and high-quality animation generation.
@@ 4.1 Datasets
We evaluate our method on three public datasets with synchronized speech and motion data. BEAT Dataset [32] contains speech, facial expressions, and body gestures from four subjects, along with annotations such as transcriptions, semantic labels, and emotional tags. Following the official protocol, we use 34-frame sequences for training/validation and 64 extended sequences (around one minute each) for testing. Motion is processed at 15 FPS using axis-angle rotation for smooth synthesis. SHOW Dataset [65] includes synchronized SMPLX [43] parameters and audio from four subjects, recorded at 30 FPS with audio sampled at 22 kHz. We use 88-frame sequences for training/validation and variable-length sequences for testing, with motion represented by SMPLX parameters. BEAT2 Dataset [33], introduced in EMAGE [33], is based on the original BEAT dataset [32], containing 76 hours of data from 30 speakers. We use the BEAT2-standard portion, splitting identities into 85%, 7.5%, and 7.5% for training, validation, and testing, and report results on BEAT2-Standard Speaker2 (Scott).

### v2

@@ Abstract
footnotetext: Equal Contribution.🖂🖂footnotetext: Corresponding authors.
@@ 1 Introduction
Audio-driven avatar expression and pose generation [18, 2, 4] is a crucial task aimed at creating lifelike digital humans that can seamlessly translate audio input into synchronized facial expressions and body poses. This task is fundamental to bridging the gap between speech and nonverbal communication, enabling avatars to convey emotions, intentions, and personality in a natural and dynamic manner. Its importance spans a wide range of fields, including metaverse applications, digital human development, gaming, and human-computer interaction [32, 27, 33].
In recent years, numerous methods have been proposed for audio-driven avatar expression and pose generation, primarily treating speech-driven facial expression and body motion synthesis as separate tasks. Facial expression generation [9, 30] focuses on mapping emotional features from speech to facial muscle movements for natural animations, while body motion synthesis [11, 8] explores correlations between speech and gestures to generate coherent full-body motions. Despite advancements, these methods often lack sufficient coordination between expressions and movements. Generative models like VQ-VAE [2], GANs [12], and diffusion models [1, 2, 31, 29] have improved synchronization and diversity, enabling unified modeling of expressions and movements [12, 7, 23, 20]. As shown in Fig. 1, recent works include Probtalk [23], which generates expressions and postures simultaneously with a unified model, DiffSHEG [7], which uses a unidirectional sequence from expression to gestures, and EMAGE [20], which incorporates body hints for better coordination. Combo [35], the most related work, combines features for expressions and movements into a joint bidirectional distribution. However, a key challenge remains: balancing coordination accuracy and computational efficiency. More specifically, synchronization of expressions and movements often incurs high computational overhead, limiting the production of fluid animations in latency-sensitive scenarios.
@@
To address this challenge, we propose AsynFusion, a framework that decouples facial expression and body gesture generation for efficient, lifelike animation. By separating head and body generation, AsynFusion enables parallel processing while maintaining coordination through shared feature interactions. This design respects the distinct dynamics of each modality and incorporates asynchronous mechanisms to improve efficiency without sacrificing quality. The model comprises three key components: (1) a dual-branch Diffusion Transformer for parallel expression-gesture generation with bidirectional interaction; (2) a cooperative synchronization module using cross-attention to capture inter-modal dependencies and enhance coherence; and (3) an asynchronous Latent Consistency Model (LCM) sampling strategy that accelerates inference while preserving motion quality, enabling real-time applications.
@@ 2.1 Speech-driven Expression Generation
Speech-driven facial animation has evolved from early rule-based methods [17, 18], which offered controllability but required manual tuning, to data-driven models [16, 24, 25] that generate more natural and speech-synchronized expressions. However, these models often suffer from pixel-level artifacts and geometric inconsistencies. Recent advances in 3D facial animation, especially transformer-based architectures like FaceFormer [10] and CodeTalker [34], have improved temporal alignment using attention mechanisms. Still, most approaches rely on deterministic mappings, limiting expression diversity. Our work builds on these developments by introducing a more expressive framework that overcomes the limitations of deterministic designs.
@@ 2.2 Speech-driven Gesture Generation
Gesture generation has similarly transitioned from rule-based systems [6, 18] to data-driven methods using MLPs, CNNs, RNNs [12, 13, 22], and Transformers [5]. Recognizing the one-to-many nature of speech-to-gesture mapping, recent works have adopted generative models like GANs [12] and diffusion models [1, 7], which offer greater motion diversity. Pioneering diffusion-based approaches such as DiffGesture [39] and DiffuseStyleGesture [36] require motion seeds, limiting their use in continuous generation. LDA [1] addressed longer sequences with translation-invariant embeddings but struggles with streaming data. Our method advances this line by enabling seed-free, continuous gesture generation in a more robust and efficient framework.
@@ 2.3 Joint Expression-Gesture Generation
Joint modeling of expressions and gestures aims to improve realism and synchronization. Habibie et al. [12] proposed a CNN-based multi-decoder system with adversarial training, though it lacked motion diversity. Yi et al. [37] used Wav2Vec [3] and VQ-VAE to decouple tasks, but tokenization constrained gesture variety. ProboTalk [23] unified expression and pose generation in one model, while EMAGE [20] used body cues to enhance expression. DiffSHEG [7] introduced a diffusion-based framework with unidirectional flow from expression to gesture, improving coordination but limiting mutual influence. Combo [35] attempted bidirectional modeling but suffered from synchronous generation inefficiencies. Our work addresses these issues with a bidirectional yet asynchronous framework, improving both interaction quality and generation efficiency without compromising diversity.
@@ 3.1 Preliminary
Diffusion Transformers Diffusion Transformers (DiT) employ a latent diffusion model (LDM) with a Transformer backbone for motion generation. Let $\mathbf{x}_{0}^{E}$, $\mathbf{x}_{0}^{G}$, and $\mathbf{x}_{0}^{M}$ denote expressions, gestures, and motion clips, respectively, aiming to model the motion distribution $p(\mathbf{x}_{0})\in\mathbb{R}^{N\times(3J+D_{exp})}$, where $N$ is the number of frames, $J$ the skeletal joints, and $D_{exp}$ the expression blend shape dimension.
@@
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$ (1)
@@
$q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$ (2)
@@
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (3)
@@
$\epsilon_{\theta}(\mathbf{x}_{t},t,c)=f_{i}\circ f_{i-1}\circ\dots\circ f_{1}(\mathbf{x}_{t},t,c),$ (4)
@@
Latent Consistency Models
@@ 3.2 The AsynFusion Framework
Dual-branch DiT Architecture. To generate coordinated facial expressions and body gestures, AsynFusion adopts a Dual-branch DiT Architecture consisting of two parallel branches: the expression branch, which captures subtle facial motions and lip synchronization, and the gesture branch, which handles broader body dynamics. Each branch uses independent transformer blocks to learn domain-specific temporal dependencies. The input to each branch includes: (1) noisy motion samples $\mathbf{z}_{t}^{E}$ or $\mathbf{z}_{t}^{G}$, obtained by adding Gaussian noise to target motions; (2) timestep embeddings $\gamma(t)$; and (3) shared audio features $\mathbf{F}_{\mathrm{aud}}$ for synchronized conditioning:
$\hat{\mathbf{z}}^{E}=\mathcal{T}_{E}(\mathbf{z}_{t}^{E},\gamma(t),\mathbf{F}_{\mathrm{aud}}),\quad\hat{\mathbf{z}}^{G}=\mathcal{T}_{G}(\mathbf{z}_{t}^{G},\gamma(t),\mathbf{F}_{\mathrm{aud}}),$ (8)
where $\mathcal{T}_{E}$ and $\mathcal{T}_{G}$ are the expression and gesture transformers. This design supports specialized learning per modality while enabling cross-branch interaction through synchronization. It also allows for asynchronous sampling to accommodate their differing temporal characteristics.
Cooperative Synchronization Module. To model the interplay between facial expressions and gestures, we introduce the Cooperative Synchronization (CoSync) module, which enables bidirectional feature exchange between branches. After each transformer block, a cross-attention-based synchronization layer captures inter-modal dependencies and enhances motion coherence.
We take gesture to expression data-flow for example, the query $\mathbf{Q}_{exp}$ is extracted by linear projection from $\mathbf{F}_{i}^{exp}$ (ii is the layer index), and the key and value $\mathbf{K}_{ges}$, $\mathbf{V}_{ges}$ are extracted from $\mathbf{F}_{i}^{ges}$ in the same way. To obtain the updated facial feature $\tilde{\mathbf{F}}_{i}^{exp}$,
@@
where MLP and LN is a MLP block and a LayerNorm, $\sqrt{d}$ is a scaling factor. This bidirectional feature exchange enables the model to capture subtle correlations between facial micro-expressions and corresponding gestural nuances, much like the natural synchronization observed in human behavior. What distinguishes the CoSync module is its ability to maintain the delicate balance between modality-specific independence and cross-modal coordination. While each branch preserves its specialized focus, the module enables them to share complementary information that enhances the overall coherence of the generated animation.
@@
The expression branch typically requires fewer sampling steps ($T_{exp}$) than the gesture branch ($T_{ges}$) due to its more constrained motion space. To support bidirectional interaction during asynchronous sampling, we introduce a dynamic feature buffer in the CoSync module. At each step, both branches store and asynchronously update their intermediate features. This allows each branch to access the latest features from the other, maintaining continuous cross-modal exchange despite differing sampling rates. As a result, the expression branch achieves fast generation for facial motions, while the gesture branch uses more steps to capture complex dynamics—balancing quality and efficiency.
@@ 3.3 Training
3.4 Long Sequence Generation
To generate arbitrary-length animations, we further integrate existing techniques in DiffSHEG [7] with our dual-branch asynchronous framework. The main challenge is ensuring smooth transitions between clips while maintaining the efficiency of asynchronous sampling.
Clip-based generation: We use a sliding window approach similar to DiffSHEG. Consecutive clips have overlapping frames to ensure smooth transitions. The starting frames of each new clip are initialized with the ending frames of the previous clip, maintaining continuity in both facial expressions and body gestures. This approach naturally fits with our asynchronous sampling mechanism, allowing each branch to sample at its own optimal rate while preserving temporal coherence.
Efficient inference: Our LCM-based method significantly reduces the number of required sampling steps compared to traditional diffusion models. The expression branch typically needs 4-6 sampling steps, while the gesture branch uses 6-8 steps—both far fewer than the 1000 steps often used in conventional models. This reduction in steps is crucial for real-time avatar animation applications. It is noteworthy that although our method employs asynchronous generation—meaning the time required to generate facial expressions and body poses for each frame may vary—the outputs are produced on a frame-by-frame basis, ensuring that the expressions and poses are aligned in each frame.
Transition refinement: To ensure smooth transitions between clips, we apply a transition refinement technique. Overlapping frames at the clip boundaries are linearly interpolated during the final sampling steps of each branch, ensuring seamless transitions while keeping expressions and gestures naturally coordinated.
@@ 4.1 Datasets
We evaluate our method on three public speech-motion datasets. BEAT Dataset [21] provides synchronized speech, facial expressions, and gestures from four subjects, along with annotations such as transcriptions, semantics, and emotions. Following the official setup, we use 34-frame clips for training/validation and 64-frame (approx. one minute) sequences for testing. Motions are represented using axis-angle rotation at 15 FPS. SHOW Dataset [37] offers synchronized SMPLX [28] parameters and audio (22 kHz) from four speakers, recorded at 30 FPS. We use 88-frame sequences for training/validation and variable-length clips for testing, with SMPLX as the motion representation.
4.2 Metrics Computation
This section outlines the evaluation metrics employed in our experimental analysis.
Fréchet Motion Distance The Fréchet Motion Distance (FMD) [38] extends the established Fréchet Gesture Distance concept, providing a reliable measure that aligns with human perceptual assessment. FMD quantifies the distributional similarity between generated and authentic motions by computing the Fréchet distance between their respective latent representations. These latent features are obtained through a specialized neural encoder trained on either the BEAT[21] or SHOW datasets[37]. The mathematical formulation is:
$FGD=|\mu_{r}-\mu_{s}|^{2}+Tr(\sigma_{r}+\sigma_{s}-2\sqrt{\sigma_{r}\sigma_{s}}),$ (16)
where $(\mu_{s},\sigma_{s})$ represent the mean and covariance statistics of the synthesized motion distribution in latent space, while $(\mu_{r},\sigma_{r})$ correspond to those of the real motion distribution. Following this framework, we define analogous metrics - Fréchet Expression Distance (FED) and Fréchet Gesture Distance (FGD) - to evaluate expression and gesture quality respectively.
Diversity metric (Div) To assess the variability of generated animations, we employ a diversity metric [4] that quantifies motion heterogeneity across batches. Given a test batch dimension B, we calculate our diversity score as:
$Div=\frac{2}{B\times(B-1)}\sum_{i=1}^{B-1}\sum_{j=i+1}^{B}|\hat{x}_{i}-\hat{x}_{j}|_{1},$ (17)
For implementation, $\hat{x}_{i}$ represents a complete motion sequence from our i-th batch generation. In our experimental protocol, we utilize a batch size B of 50 samples to ensure robust diversity assessment.
Beat Alignment (BA) To evaluate temporal coherence between audio and generated movements, we implement the Beat Alignment (BA) metric. This assessment tool examines the temporal correlation by quantifying the proximity between motion-derived beats and their audio counterparts. The mathematical representation is:
$BA=\frac{1}{n}\sum_{i=1}^{n}\exp{-\frac{\min_{\forall b_{j}^{a}\in B^{a}}|b_{i}^{m}-b_{j}^{a}|^{2}}{2\sigma^{2}}},$ (18)
In this formulation, $B^{a}$ represents the set of detected audio beats $\{b_{i}^{m}\}$, while $B^{m}$ encompasses the extracted motion beats $\{b_{i}^{m}\}$. Our implementation adheres to the standardized beat detection and alignment procedures established in the BEAT [21] and TalkSHOW [37].
4.3 Implementation Details
Experiments are conducted on two NVIDIA A100 (40G) GPUs. On BEAT, we train for 1,000 epochs with a batch size of 1,600. On SHOW, due to longer sequences and higher frame rates, we train for 1,600 epochs with a batch size of 700. We compare AsynFusion with recent state-of-the-art methods, focusing on DiffSHEG [7] and Combo [35] for joint generation. All models use axis-angle rotation and are conditioned on audio and speaker identity. For gesture-only models (DiffGesture [39], DiffuseStyleGesture [36], LDA [1]), we train separate facial models for fair comparison. Although our focus is on upper-body motion, AsynFusion supports full-body synthesis. We pay particular attention to comparing our bidirectional interaction and asynchronous sampling with the unidirectional design of DiffSHEG [7].